# DrivAerNet: A Parametric Car Dataset for Data-Driven Aerodynamic Design and Graph-Based Drag Prediction Аннотация

### v1

@@ 4 Dynamic Graph Convolutional Neural Network for Regression
In this work, we extend the Dynamic Graph Convolutional Neural Network (DGCNN) framework [42], traditionally associated with PointNet [27] and graph CNN methodologies, to address regression tasks, marking a significant departure from its conventional applications in classification. Our contribution lies in adapting the DGCNN architecture to predict continuous values, specifically focusing on aerodynamic coefficients critical in fluid dynamics and engineering design. Leveraging the spatial encoding capabilities of PointNet and the relational inferences provided by graph CNNs, our proposed RegDGCNN model (as shown in Figure 10) aims to capture the complex interactions of fluid flow around objects, offering a novel method for accurate estimation of crucial aerodynamic parameters. This approach harnesses local geometric structures by constructing a local neighborhood graph and applying convolution-like operations on the edges linking pairs of neighboring points, aligning with graph neural network principles. The technique, termed edge convolution (EdgeConv), is shown to exhibit properties that bridge translation invariance and non-locality. Uniquely, unlike in standard graph CNNs, the graph of RegDGCNN is not static but is dynamically updated after each layer of the network, allowing the graph structure to adapt to evolving feature spaces.
We start by initializing the graph with node features , along with the parameters for the EdgeConv layers and the Fully Connected (FC) layers . A distinctive feature of the RegDGCNN is its dynamic graph construction within each EdgeConv layer, where the -nearest neighbors of each node are identified based on the Euclidean distance in the feature space, thereby adaptively updating the graph’s connectivity to reflect the most significant local structures. The EdgeConv operation is defined as:
enhances node features by aggregating information from these neighbors using a shared Multi-Layer Perceptron (MLP), which processes both the individual node features and their differences with adjacent nodes, capturing the local geometric context effectively.
@@
Here, max pooling is employed to encapsulate the graph’s holistic information. This global feature vector is subsequently processed through several FC layers, with the inclusion of non-linear activation functions like ReLU and dropout to introduce non-linearity and prevent overfitting, respectively. The architecture culminates in an output layer, designed to suit the specific task at hand, for instance, employing a linear activation for regression tasks.
The model’s performance is quantified by calculating the loss between its predicted outputs and the ground truth drag values using the mean squared error (MSE), with the backpropagation algorithm adjusting the model parameters and through optimization algorithms such as Adam [21] to minimize this loss. This iterative refinement process highlights the RegDGCNN’s capability to dynamically leverage and integrate hierarchical features from graph-structured data.
@@ 4.1 Implementation Details
We constructed the graph for the RegDGCNN using the -nearest neighbors algorithm, with set to 40. This parameter was critical in defining the local neighborhood over which convolutional operations were performed. The RegDGCNN model was instantiated with specific parameters to accommodate the nature of the regression task. The EdgeConv layers were configured with channels of sizes {256, 512, 512, 1024}, and the following MLP layers were {128, 64, 32, 16}. Lastly, the embedding dimension of the network was set at 512, providing a high-dimensional space to capture the complex features necessary for the regression tasks at hand. Our RegDGCNN model is entirely differentiable and can seamlessly integrate with 3D generative AI applications for enhancing design optimization.
@@
The experiments were conducted using the PyTorch framework. The models were trained using a batch size of 32, with the number of points per input set to 5000. Training was distributed across 4 NVIDIA A100 80GB GPUs, utilizing data parallelism to enhance computational efficiency. The network’s learning rate was initially set to 0.001, and a learning rate scheduler was employed to reduce the rate upon plateauing of the validation loss, specifically using the ReduceLROnPlateau scheduler with a patience of 10 epochs and a reduction factor of 0.1. This approach helped in fine-tuning the models by adjusting the learning rate in response to the performance on the validation set. The models were trained for a total of 100 epochs, ensuring sufficient learning while preventing overfitting. For optimization, we used the Adam optimizer [21] due to its adaptive learning rate capabilities.
Inference Time:
The RegDGCNN model, with its compact size of roughly 3 million parameters and a storage requirement of about 10MB, performs drag estimation for a car design with 540k mesh faces in 1.2 seconds on 4 A100 80GB GPUs, a significant efficiency gain compared to the 2.3 hours taken by a standard CFD simulation on 128 CPU cores with 4 A100 80GB GPUs.
@@ 5 Surrogate Modeling of Aerodynamic Drag
In this section, we evaluate our RegDGCNN model on two aerodynamic datasets, DrivAerNet and ShapeNet, highlight the impact of larger training volumes on model performance, and investigate the model’s learned features.
@@ 5.1 DrivAerNet: Aerodynamic Drag Prediction of High-Resolution Meshes
5.2 ShapeNet: Aerodynamic Drag Prediction of Arbitrary Vehicle Shapes
@@
In Table 3, we compare the performance of two models: the attn-ResNeXt model from the study by [36], which implements a self-attention mechanism to boost the understanding of interactions among various regions of an image. It uses 2D depth/normal renderings as input and has approximately 2 billion parameters, achieving an score of 0.84; our proposed RegDGCNN model, which directly processes 3D mesh data, significantly reduces the number of parameters to 3 million, and achieves a superior score of 0.87. This comparison underscores the efficiency and effectiveness of our model in aerodynamic drag prediction tasks.
@@ 5.3 Effect of Training Dataset Size
For both ShapeNet444The work expanded their dataset to 9,896 variations through resizing and flipping augmentations, but considered only 2,474 unique car designs as independent samples for their surrogate model training to prevent data leakage across different sets. and DrivAerNet datasets, we first allocated 70 for training, and 15 each for validation and testing. Subsequently, we experimented with training subsets at 20, 40, 60, 80, and 100 of the training portion. The ShapeNet subsets ranged from 1270 to 6352 samples. Meanwhile, for the DrivAerNet dataset, the corresponding sample sizes were 560, 1120, 1680, 2240, and 2800 samples.
Figure 13 reveals a clear trend where the average relative error in drag coefficient predictions decreases as the percentage of the dataset used for training increases. This trend is consistent for both datasets, underscoring the common machine learning principle that more training data generally leads to better model performance. The improved performance of the DrivAerNet Dataset across all sizes of training data highlights the critical role of bigger datasets in machine learning models for aerodynamics and further establishes the value of DrivAerNet dataset, which is significantly larger than previous open-source datasets.
@@
ShapeNet models cars were modeled as single-bodied entities, omitting crucial details like wheels and underbodies that are vital for accurate aerodynamic modeling.
All drag values in ShapeNet were computed using a single reference area, which does not account for the significant variation in frontal projected areas of different car designs.
@@
This analysis serves to demonstrate the generalization capabilities of our model, emphasizing the objective of developing models that effectively generalize to out-of-domain distributions.
5.4 Features Learned
To further assess the model’s performance, we analyze the features learned in the intermediate layers following the edge convolution operations. Figure 14 illustrates the distribution of feature importance across an upsampled point cloud of a car sample from DrivAerNet, with color coding from light yellow (low importance) to dark red (high importance). Initially, RegDGCNN zeroes in on the car’s frontal and rear area, which are critical in shaping aerodynamic performance. This focus is notably pertinent for aerodynamic design, as the frontal area has a substantial effect on pressure drag and the rear area is particularly relevant due to its role in airflow separation and wake region formation. As the model delves deeper through its layers, it starts to recognize more complex geometric details. Conversely, areas such as the roof and windows are less crucial for drag, underscoring the model’s adeptness at identifying areas with a more significant aerodynamic influence.
@@ 6 Limitations and Future Work
This section discusses the limitations of our study. Despite careful selection to ensure a balance between detail and computational efficiency, the model’s parameterization faces an inherent limitation. This stems from the trade-off between the compactness of the representation and the flexibility needed to capture a broad range of aerodynamic phenomena. Consequently, while our approach offers valuable insights for many applications, it may not fully encompass all aerodynamic variabilities pertinent to automotive engineering. The dataset comprises 4,000 instances, which, while significant, might not fully capture the broad spectrum of real-world automotive designs. Additionally, our focus was primarily on drag prediction; however, we plan to extend the application of RegDGCNN to include surface field predictions in future work. While our dataset is large and of high-fidelity, it is important to acknowledge that we are still in the early stages of approaching the scale and foundational impact seen in AI fields like image processing and natural language processing, where large datasets are a norm.
One of the key challenges in applying graph-based methods like RegDGCNN is the significant GPU memory requirements. This is due to the need to compute all pairwise distances between points, which can be highly memory-intensive. Moreover, the non-uniform density of point clouds introduces additional complexities; a fixed -nearest neighbors approach may not be suitable for areas with varying densities of points.
Another limitation is that, in its current form, RegDGCNN does not reduce the number of points during the forward pass for large-scale point clouds, leading to high computational demands and potentially limiting the model’s scalability to even larger datasets. Addressing these challenges will be crucial for advancing the capabilities and applications of graph-based neural networks in processing complex aerodynamic data.
@@ 7 Conclusion
In our conclusion, we highlight the distinct advantages of DrivAerNet, which, by focusing on detailed geometric modifications, outperforms broader datasets such as those referenced in [36, 22, 31], especially in the context of real-world aerodynamic design applications. Additionally, the compact RegDGCNN model, with 3 million parameters and 10MB size, efficiently estimates drag in just 1.2 seconds for industry-standard designs with 540k mesh faces, significantly outpacing traditional CFD simulations. Moreover, our RegDGCNN model, in particular, showcases superior performance by directly processing 3D meshes, thereby eliminating the need for 2D image rendering or the generation of Signed Distance Functions (SDF), which simplifies the preprocessing stages and increases the model’s accessibility. Importantly, the RegDGCNN model’s ability to deliver precise drag predictions without requiring water-tight meshes highlights its adaptability and effectiveness in leveraging real-world data. Through the expansion of the DrivAerNet dataset from 560 to 2800 samples, we achieved a remarkable reduction in error by approximately 75. Similarly, on the dataset from [36], increasing the training samples from 1270 to 6352 led to a 56 decrease in error, underscoring the substantial impact of dataset scale on enhancing the performance of deep learning models in aerodynamic studies. The inclusion of specific parametric alterations (50 geometric parameters) within our DrivAerNet dataset has significantly improved model learning, resulting in a notable increase in predictive accuracy, essential for the refinement of aerodynamic designs. This emphasizes the critical role of large detailed, high-fidelity datasets in crafting models capable of adeptly handling the complexities inherent in aerodynamic surrogate modeling.

### v2

@@ 5 Dynamic Graph Convolutional Neural Network for Regression
In this work, we extend the Dynamic Graph Convolutional Neural Network (DGCNN) framework [50], traditionally associated with PointNet [35] and graph CNN methodologies, to address regression tasks, marking a significant departure from its conventional applications in classification. Our contribution lies in adapting the DGCNN architecture to predict continuous values, specifically focusing on aerodynamic coefficients critical in fluid dynamics and engineering design. Leveraging the spatial encoding capabilities of PointNet and the relational inferences provided by graph CNNs, our proposed RegDGCNN model (as shown in Figure 10) aims to capture the complex interactions of fluid flow around objects, offering a novel method for accurate estimation of crucial aerodynamic parameters. This approach harnesses local geometric structures by constructing a local neighborhood graph and applying convolution-like operations on the edges linking pairs of neighboring points, aligning with graph neural network principles. The technique, termed edge convolution (EdgeConv), is shown to exhibit properties that bridge translation invariance and non-locality. Uniquely, unlike in standard graph CNNs, the graph of RegDGCNN is not static but is dynamically updated after each layer of the network, allowing the graph structure to adapt to evolving feature spaces. This dynamic graph construction results in better performance for 3D geometric data while maintaining computational efficiency during both training and inference, as the graph is built on top of the sampled point clouds. Moreover, DGCNN is well-suited for handling unstructured point cloud data, which aligns well with the nature of the mesh format in our dataset.
We start by initializing a graph with node features , along with the parameters for the EdgeConv layers and the Fully Connected (FC) layers . A distinctive feature of the RegDGCNN is its dynamic graph construction within each EdgeConv layer, where the -nearest neighbors of each node are identified based on the Euclidean distance in the feature space, thereby adaptively updating the graph’s connectivity to reflect the most significant local structures. The EdgeConv operation, defined as:
enhances node features by aggregating information from these neighbors using a shared Multi-Layer Perceptron (MLP), which processes both the individual node features and their differences with adjacent nodes, capturing the local geometric context effectively. For the EdgeConv layers, we use same convolution with a kernel size of 1, meaning the input and output dimensions remain unchanged throughout each EdgeConv operation. This helps maintain the feature space dimensions while effectively capturing local geometric features through the dynamic graph structure.
@@
Here, max pooling is employed to encapsulate the graph’s holistic information. This global feature vector is subsequently processed through several FC layers, with the inclusion of non-linear activation functions like LeakyReLU and dropout to introduce non-linearity and prevent overfitting, respectively. The architecture culminates in an output layer, designed to suit the specific task at hand, for instance, employing a linear activation for regression tasks.
The model’s performance is quantified by calculating the loss between its predicted outputs and the ground truth drag values using the mean squared error (MSE), with the backpropagation algorithm adjusting the model parameters and through optimization algorithms such as Adam [26] to minimize this loss. This iterative refinement process highlights RegDGCNN’s capability to dynamically leverage and integrate hierarchical features from graph-structured data.
@@ 5.1 Implementation Details
We constructed the graph for RegDGCNN using the -nearest neighbors algorithm, with set to 40. This parameter was critical in defining the local neighborhood over which convolutional operations were performed. The RegDGCNN model was instantiated with specific parameters to accommodate the nature of the regression task. The EdgeConv layers were configured with channels of sizes {256, 512, 512, 1024}, and the following MLP layers were {128, 64, 32, 16}. Lastly, the embedding dimension of the network was set at 512, providing a high-dimensional space to capture the complex features necessary for the regression tasks at hand. Our RegDGCNN model is entirely differentiable and can seamlessly integrate with 3D generative AI applications for enhancing design optimization.
@@
The experiments were conducted using the PyTorch framework [32]. The model was trained with a batch size of 32 and each input containing 5,000 points. The training was distributed across four NVIDIA A100 80GB GPUs using data parallelism for improved computational efficiency, and the process took approximately 3.2 hours to complete. The network’s learning rate was initially set to 0.001, and a learning rate scheduler was employed to reduce the rate upon plateauing of the validation loss, specifically using the ReduceLROnPlateau scheduler with a patience of 10 epochs and a reduction factor of 0.1. This approach helped in fine-tuning the model by adjusting the learning rate in response to the performance on the validation set. The model was trained for a total of 100 epochs, ensuring sufficient learning while preventing overfitting. For optimization, we used the Adam optimizer [26] due to its adaptive learning rate capabilities.
Evaluation Metrics
Various loss functions are used to evaluate the accuracy of models in predicting the aerodynamic drag of different cars. The primary metrics are described below:
Mean Squared Error (MSE): This metric calculates the average of the squared discrepancies, which represent the differences between the actual drag coefficients obtained from CFD simulations and those predicted by the surrogate models. It is particularly responsive to large deviations in predictions.
(16) -
Coefficient of Determination ( Score): The Score quantifies the fraction of variance in the observed drag coefficients that can be accounted for by the model’s predictions. An value of 1 indicates a perfect prediction.
@@ 6 Surrogate Modeling of Aerodynamic Drag
In this section, we evaluate our RegDGCNN model on two aerodynamic datasets, DrivAerNet and ShapeNet, and highlight the impact of larger training volumes on model performance.
@@ 6.1 DrivAerNet: Aerodynamic Drag Prediction of High-Resolution Meshes
6.2 ShapeNet: Aerodynamic Drag Prediction of Arbitrary Car Shapes
@@
In Table 4, we compare the performance of two models: the attn-ResNeXt model from the study by [44], which implements a self-attention mechanism to boost the understanding of interactions among various regions of an image. It uses 2D depth/normal renderings as input and has approximately 2 billion parameters, achieving an score of 0.84; our proposed RegDGCNN model, which directly processes 3D mesh data, significantly reduces the number of parameters to 3 million, and achieves a superior score of 0.87. The superior performance of RegDGCNN compared to attn-ResNeXt can be attributed to the fact that the model uses 1000x fewer parameters and works directly on the 3D geometry rather than 2D rendered depth/normal maps, which allows it to learn more features and helps to avoid overfitting. This comparison underscores the efficiency and effectiveness of our model in aerodynamic drag prediction tasks.
@@ 6.3 Effect of Training Dataset Size
For both ShapeNet222The work expanded their dataset to 9,896 variations through resizing and flipping augmentations but considered only 2,474 unique car designs as independent samples for their surrogate model training to prevent data leakage across different sets. and DrivAerNet datasets, we first allocated 70 for training, and 15 each for validation and testing. Subsequently, we experimented with training subsets at 20, 40, 60, 80, and 100 of the training portion. The ShapeNet subsets ranged from 1270 to 6352 samples. Meanwhile, for the DrivAerNet dataset, the corresponding sample sizes were 560, 1120, 1680, 2240, and 2800 samples.
Figure 13 reveals a clear trend where the average relative error in drag coefficient predictions decreases as the percentage of the dataset used for training increases. This trend is consistent for both datasets, underscoring the common machine learning principle that more training data generally leads to better model performance. The improved performance of the DrivAerNet Dataset across all sizes of training data highlights the critical role of bigger datasets in machine learning models for aerodynamics and further establishes the value of the DrivAerNet dataset, which is significantly larger than previous open-source datasets.
@@
ShapeNet cars were modeled as single-bodied entities, omitting crucial details like wheels and underbodies that are vital for accurate aerodynamic modeling.
@@
This analysis serves to demonstrate the generalization capabilities of our model, emphasizing the objective of developing models that effectively generalize to out-of-distribution designs.
@@ 7 Limitations and Future Work
This section discusses the limitations of our study. Despite careful selection to ensure a balance between detail and computational efficiency, the model’s parameterization faces an inherent limitation. This stems from the trade-off between the compactness of the representation and the flexibility needed to capture a broad range of aerodynamic phenomena. Consequently, while our approach offers valuable insights for many applications, it may not fully encompass all aerodynamic variabilities pertinent to automotive engineering. The dataset comprises 4,000 instances, which, while significant, might not fully capture the broad spectrum of real-world automotive designs. A feature importance analysis should be conducted on the parametric data to efficiently determine the impact of various parameters on aerodynamic performance. This approach will facilitate more efficient sampling and guide the prioritization of parameters in future dataset expansions. While our dataset is large and of high-fidelity, it is important to acknowledge that we are still in the early stages of approaching the scale and foundational impact seen in AI fields like image processing and natural language processing, where large datasets are a norm.
One of the key challenges in applying graph-based methods like RegDGCNN is the significant GPU memory requirements. This is due to the need to compute all pairwise distances between points, which can be highly memory-intensive. Moreover, the non-uniform density of point clouds introduces additional complexities; a fixed -nearest neighbors approach may not be suitable for areas with varying densities of nodes. Another limitation is that, in its current form, RegDGCNN does not reduce the number of points during the forward pass for large-scale point clouds, leading to high computational demands and potentially limiting the model’s scalability to even larger datasets. Additionally, our focus was primarily on drag prediction; however, we plan to extend the application of RegDGCNN to include surface field predictions in future work. Addressing these challenges will be crucial for advancing the capabilities and applications of graph-based neural networks in processing complex aerodynamic data.
@@ 8 Conclusion
In our conclusion, we highlight the distinct advantages of DrivAerNet, which, by focusing on detailed geometric modifications, outperforms broader datasets such as those referenced in [44, 28, 39], especially in the context of real-world aerodynamic design applications. Additionally, the compact RegDGCNN model, with 3 million parameters and a 10MB size, efficiently estimates drag in just 1.2 seconds for industry-standard designs with 540k mesh faces, significantly outpacing traditional CFD simulations. Moreover, our RegDGCNN model, in particular, showcases superior performance by directly processing 3D meshes, thereby eliminating the need for 2D image rendering or the generation of Signed Distance Functions (SDF), which simplifies the preprocessing stages and increases the model’s accessibility. Importantly, the RegDGCNN model’s ability to deliver precise drag predictions without requiring water-tight meshes highlights its adaptability and effectiveness in leveraging real-world data. Through the expansion of the DrivAerNet dataset from 560 to 2800 samples, we achieved a remarkable reduction in error by approximately 75. Similarly, on the dataset from [44], increasing the training samples from 1270 to 6352 led to a 56 decrease in error, underscoring the substantial impact of dataset scale on enhancing the performance of deep learning models in aerodynamic studies. The DrivAerNet dataset, which is generated through the morphing of a baseline parametric model to produce high-quality industry-standard designs, proves to be a better approach compared to the designs found in ShapeNet. This emphasizes the critical role of large, detailed, high-fidelity datasets in crafting models capable of adeptly handling the complexities inherent in aerodynamic surrogate modeling.
9 Dataset Name
Name: DrivAerNet
10 Motivation
10.1 Why was the dataset created?
The DrivAerNet dataset was developed to provide a benchmark for training deep learning models in aerodynamic car design. It was created to address the lack of publicly available datasets that include industry-standard parametric car geometries and high-fidelity computational fluid dynamics (CFD) simulations, enabling researchers to develop and test machine learning models for engineering applications.
10.2 Who created the dataset?
DrivAerNet was created by the Design Computation and Digital Engineering (DeCoDE) Lab at the Massachusetts Institute of Technology (MIT) in collaboration with the 3D AI Lab at the Technical University of Munich (TUM).
10.3 Who funded the creation of the dataset?
The dataset creation was primarily supported by computational resources provided by MIT. Mohamed Elrefaie was supported by MIT DeCoDE lab with additional support from TUM and the DAAD IFI Fellowship.
11 Composition
11.1 What do the instances that comprise the dataset represent?
Each instance in DrivAerNet represents a parametric 3D car geometry along with high-fidelity CFD simulation results. The dataset includes 3D fields for velocity, pressure, and wall-shear stresses, as well as aerodynamic coefficients for various car configurations.
11.2 How many instances are there in total?
DrivAerNet consists of 4,000 full-domain 3D CFD simulations.
11.3 What data does each instance consist of?
Each instance contains car geometry files in STL format, CFD simulation results in VTK format, aerodynamic coefficients, and tabular parametric design data. Additional annotations and point cloud representations are also included for machine learning applications.
11.4 Is there any missing data? If so, how much?
No.
11.5 How is the data associated with each instance organized?
Each design is stored in a structured folder system, with separate files for geometry and simulation results ensuring easy access and reproducibility.
12 Collection Process
12.1 How was the data collected?
The dataset was generated using parametric car designs that were simulated in OpenFOAM, with mesh generation performed using SnappyHexMesh. Python scripts automated the data generation and extraction process.
12.2 Who was involved in the data collection process?
The dataset was developed by researchers from MIT and TUM, utilizing automated pipelines to generate, process, and validate the simulations.
12.3 Over what timeframe was the data collected?
The dataset was created over the course of 2023 and 2024.
12.4 How was it decided which data to collect and which to exclude?
Only geometries that passed validation checks for CFD simulations were included, ensuring accurate and meaningful aerodynamic analysis.
13 Preprocessing, Cleaning, and Labeling
13.1 Was any preprocessing, cleaning, or labeling done?
Yes, preprocessing was applied to remove geometries with non-manifold edges or intersecting surfaces. The dataset also includes labeled aerodynamic coefficients and part annotations for machine learning applications.
13.2 How was this done and by whom?
The preprocessing was automated using Python scripts developed by the MIT and TUM research teams.
14 Uses
14.1 For what tasks in Engineering Design is the dataset suitable?
DrivAerNet is suitable for aerodynamic shape optimization, turbulence modeling, generative AI for car design, reduced-order modeling, and surrogate modeling of aerodynamic forces.
14.2 Has the dataset been used for any tasks already?
Yes, DrivAerNet has been utilized for aerodynamic surrogate modeling of the aerodynamic drag.
14.3 What performance metrics are relevant for assessing tasks using this dataset?
Relevant metrics include drag and lift coefficient prediction accuracy, R-squared for regression models, and Chamfer distance for design similarity tasks.
15 Distribution
15.1 How is the dataset distributed?
DrivAerNet is publicly available on GitHub https://github.com/Mohamedelrefaie/DrivAerNet and will also be hosted on the Harvard Dataverse https://dataverse.harvard.edu/dataverse/DrivAerNet for long-term availability.
15.2 Are there any restrictions or licenses on its use?
DrivAerNet is released under the Creative Commons Attribution-NonCommercial 4.0 International License (CC BY-NC 4.0) and is exclusively for non-commercial research and educational purposes.
15.3 What is the duration of availability for the dataset as provided by the authors?
The dataset will remain publicly available indefinitely, with updates provided as needed.
16 Maintenance
16.1 Who is responsible for dataset maintenance?
The dataset is maintained by the Harvard Dataverse team and the authors.
16.2 How can individuals submit corrections or updates?
Corrections and updates can be submitted via GitHub issues https://github.com/Mohamedelrefaie/DrivAerNet/issues and pull requests https://github.com/Mohamedelrefaie/DrivAerNet/pulls.
16.3 Is there a versioning system in place?
Yes, the dataset follows a versioning system to track updates and modifications, ensuring reproducibility.
Nomenclature
Roman and Greek Letters
= | Velocity components in the , , and directions, respectively (m/s) | |
= | freestream velocity (m/s) | |
= | Pressure (Pa) | |
= | Fluid density (kg/m3) | |
= | Dynamic viscosity (Pa·s or kg/(m·s)) | |
= | Body force components in the , , and directions (N) | |
= | Turbulence kinetic energy (m2/s2) | |
= | Specific dissipation rate (1/s) | |
= | Generation of turbulence kinetic energy due to mean velocity gradients | |
= | Diffusivities of turbulence kinetic energy and specific dissipation rate, respectively | |
= | Source terms for turbulence kinetic energy and specific dissipation rate, respectively | |
= | Magnitude of the mean rate-of-strain tensor | |
= | Kinematic viscosity, (m2/s) | |
= | Turbulence viscosity (m2/s) | |
= | Model constants | |
= | Reynolds stress tensor (Pa) |
Dimensionless Groups
Re | = | Reynolds number, |
= | Drag coefficient | |
= | Lift coefficient | |
= | Rear lift coefficient | |
= | Front lift coefficient | |
= | Pressure coefficient, |
Machine Learning and Graph Terms
= | Graph representing nodes and edges | |
= | Node features of the graph | |
= | Edge features between node and node | |
= | Parameters of the EdgeConv layers | |
= | Parameters of the Fully Connected layers | |
= | Neighborhood of node | |
= | Max pooling operation for global feature aggregation |
Chamfer Distance Terms
= | Chamfer distance between point sets and | |
= | Cardinality (number of points) in point sets and , respectively | |
= | Points in point sets and , respectively | |
= | Squared Euclidean distance between point and point |
Evaluation Metrics
= | Actual drag coefficient obtained from CFD simulations | |
= | Predicted drag coefficient from the surrogate model | |
= | Mean drag coefficient | |
= | Number of samples | |
MSE | = | Mean Squared Error, |
= | Coefficient of Determination, |
Subscripts and Superscripts
= | Indices used in Einstein summation notation, representing Cartesian coordinates | |
= | Denotes Reynolds-averaged quantities | |
= | Denotes fluctuating quantities |
Miscellaneous
= | Blending function in the SST model | |
= | Model constants used in the SST model |
Acronyms
SDF | = | Signed Distance Fields |
CFD | = | Computational Fluid Dynamics |
HPC | = | High-Performance Computing |
RegDGCNN | = | Dynamic Graph Convolutional Neural Network for Regression |
MLP | = | Multi-Layer Perceptron |
FC | = | Fully Connected |
RANS | = | Reynolds-Averaged Navier-Stokes |
URANS | = | Unsteady Reynolds-Averaged Navier-Stokes |
LBM | = | Lattice Boltzmann Method |
LES | = | Large Eddy Simulation |
SST | = | Shear Stress Transport |
PDEs | = | Partial Differential Equations |
DNN | = | Deep Neural Networks |
PCA | = | Principal Component Analysis |
GINO | = | Geometry-Informed Neural Operator |
GeoCA | = | Geometry-Guided Conditional Adaptation |
KNN | = | K-Nearest Neighbors |
OpenFOAM | = | Open Source Field Operation and Manipulation |
SHM | = | SnappyHexMesh |
STL | = | Stereolithography |
VTK | = | Visualization Toolkit |
CSV | = | Comma-Separated Values |
KDE | = | Kernel Density Estimation |
ESE | = | Enhanced Stochastic Evolutionary |